Prashnikaप्रश्निका
Code
About
‹ Back to the paper
Mathematics20262 marksNumericalProbabilityTool: Venn diagram generator and visualiser

Solve the following.

If A and B are two independent events such that P(A̅ ∪̅ B̅) = 1/3 and P(A̅) = 2/3, calculate P(B).

If $A$ and $B$ are two independent events such that $P(\overline{A \cup B}) = \frac{1}{3}$ and $P(\bar{A}) = \frac{2}{3}$, calculate $P(B)$.

Answer

Answer

$P(A \cup B) = 1 - \frac{1}{3} = \frac{2}{3}$, $P(A) = 1 - \frac{2}{3} = \frac{1}{3}$. For independent events, $P(A \cup B) = P(A) + P(B) - P(A)P(B)$. $\frac{2}{3} = \frac{1}{3} + P(B) - \frac{1}{3}P(B)$ $\frac{2}{3}P(B) = \frac{1}{3}$ Hence $P(B) = \frac{1}{2}$.

Final answer: $\frac{1}{2}$